OpenStudy (anonymous):

Evaluate [f(x+h)-f(x)]/h if (x^2) -2x

jimthompson5910 (jim_thompson5910):

\[\large f(x)=x^2-2x\] \[\large f(x+h)=(x+h)^2-2(x+h)\] \[\large f(x+h)=x^2+2xh+h^2-2x-2h\] ------------------------------------------ \[\large \frac{f(x+h)-f(x)}{h}\] \[\large \frac{(x^2+2xh+h^2-2x-2h)-(x^2-2x)}{h}\] \[\large \frac{x^2+2xh+h^2-2x-2h-x^2+2x}{h}\] \[\large \frac{2xh+h^2-2h}{h}\] \[\large \frac{h(2x+h-2)}{h}\] \[\large 2x+h-2\] \[\large 2x-2+h\] So \[\large \frac{f(x+h)-f(x)}{h}=2x-2+h\]

OpenStudy (anonymous):

Thanks a bunch Bro! WWWYKI